# Efficient Semi-Supervised Adversarial Training via Latent Clustering-Based Data Reduction

### v2

@@ III-A Prioritize Boundary-Adjacent Unlabeled Data
Algorithm 1 Prediction Confidence-Based Selection
4:  Initialize $\mathcal{A}_{u}$ as empty set
6:  Sort $\mathcal{S}_{u}$ by ascending order of $\mathrm{Conf}(\bm{x})$
III-B Strategic Selection via Latent Clustering
PCS. To identify a small, critical subset of boundary-adjacent data points while accounting for computational efficiency and scalability, we first propose a straightforward approach, coined prediction confidence-based selection (PCS), which makes use of the intermediate model $f_{\hat{\theta}}$ to compute a prediction confidence score for each unlabeled data point. The pseudocode for such a selection scheme is depicted in Algorithm 1. Initially, all the data points in $\mathcal{S}_{u}$ are sorted by their prediction confidence $\mathrm{Conf}(\cdot)$, with those exhibiting the lowest confidence scores being prioritized. The underlying assumption is that data points with low confidence scores are more likely to lie near the model’s decision boundary, making them ideal candidates for our selection. Note that the parameter $\beta\in[0,1]$ is introduced to balance the ratio between boundary and non-boundary points to avoid overfitting, which is used in all our proposed selection schemes. The reason for involving such a trade-off parameter will be further discussed in Section V-C.
@@ III-B Strategic Selection via Latent Clustering
In particular, we explore two classical clustering techniques for their unique benefits in identifying boundary points in the LCS framework. Algorithm 2 presents the pseudocode of the two proposed LCS methods. In particular, LCS with k-means (LCS-KM) generates latent representations of unlabeled data and clusters them based on Euclidean distances to the centroids, prioritizing data points that are equidistant from multiple centroids to capture local geometric structures around decision boundaries effectively. On the other hand, LCS with Gaussian mixture models (LCS-GMM) fits the latent representations to Gaussian mixture models, using posterior probabilities across multiple fitted Gaussians to identify points that are likely near decision boundaries. Compared to PCS, both LCS-KM and LCS-GMM leverage latent space clustering to provide a more accurate characterization of boundary vulnerabilities.
Algorithm 2 Latent Clustering-Based Selection
4:  Initialize $\mathcal{A}_{u}$ as empty set
6:  if using LCS-KM then
10:  else if using LCS-GMM then
11:   Fit a GMM with $k$ components to $\{\bm{z}\}_{\bm{x}\in\mathcal{S}_{u}}$
12:   Compute GMM posterior probability $\bm{p}$ for each $\bm{z}$
15:  end if
@@
where $\pi_{j}$ is the mixing coefficient for the $j$-th Gaussian component, and $\mathcal{N}(\bm{z}\mid\bm{\mu}_{j},\mathbf{\Sigma}_{j})$ denotes the Gaussian distribution with mean $\bm{\mu}_{j}$ and covariance $\Sigma_{j}$. We focus on data points with similar top posterior probabilities, suggesting that they are near the boundary between Gaussian distributions. Specifically, for each $\bm{z}$, we calculate the difference between the highest and the second-highest posterior probabilities: $\Delta p=|p_{1}(\bm{z})-p_{2}(\bm{z})|$, where $p_{1}(\bm{z})$ and $p_{2}(\bm{z})$ are the highest and second-highest posterior probabilities, respectively. Data points with the smallest $\Delta\gamma$ values are selected, as they are more likely to be near decision boundaries. Finally, the top $\alpha|\mathcal{S}_{u}|$ unlabeled data with the smallest $\Delta p$ values are used to form the selected subset. As will be illustrated in Section IV, the robustness performance of SSAT can be largely maintained when using LCS-KM and LCS-GMM to select a small subset of unlabeled data (e.g., $\alpha=10\%$ or $\alpha=20\%$). More detailed discussions of the difference between these data selection schemes are provided in Section V-B, where we visualize the selected unlabeled data in a two-dimensional latent space.
Algorithm 3 SSAT with Guided DDPM Fine-Tuning
3:  Initialize $\mathcal{G}_{u}$ as empty set
5:  Initialize $\theta$ as $\theta_{\mathrm{pre}}$
6:  for $s=1,2,\ldots,S$ do
10:   if using PCG then
12:   else if using LCG-KM then
14:   else if using LCG-GMM then
16:   end if
18:  end for
III-C Guided Diffusion via Latent Clustering
@@
While our selection-based strategy can significantly reduce the unlabeled data set size and lower the overall training time, it has a critical drawback–when SSAT employs a generative approach, all our previous methods require generating the full synthetic unlabeled dataset as a prerequisite before selection, which is inefficient in terms of both generation time and storage requirement. As shown in Table II, generating 1M CIFAR-10-like synthetic images via DDPM requires around $4$ hours on a single A100 GPU. Note that the best current SSAT method documented in the RobustBench leaderboard [17] utilizes a few hundred million synthetically generated images, suggesting the total generation time can be a significant computational overhead. Therefore, a natural question arises: whether pre-generating synthetic data and then applying strategic selection schemes can be further improved. In this section, we propose an alternative generative pipeline, detailed in Algorithm 3, which answers the above question affirmatively. In particular, we fine-tune a pretrained DDPM using novel guidance losses for more efficient unlabeled data generation based on the insights from previous sections. In this way, we directly generate a small, critical set of boundary-adjacent data points, thereby avoiding redundant data generation and further reducing the computation overhead.
@@
DDPM. Diffusion model resembles a milestone in the history of generative modeling, which is the current state-of-the-art for high-quality image generation and synthesis. A denoising diffusion probabilistic model (DDPM) is a deep neural network trained to learn how to step-by-step reverse the forward process of adding Gaussian noise to an image [23]. By gradually removing Gaussian noise through the reverse process, DDPM can thus generate real-looking images from pure noise. Let $q_{0}$ be a real image distribution over $\mathbb{R}^{d}$ and $\mathcal{N}(\bm{0},\mathbf{I})$ be the noise distribution. The forward process defines how to map images $\bm{x}_{0}\sim q_{0}$ to Gaussian noises $\bm{x}_{T}\sim\mathcal{N}(\bm{0},\mathbf{I})$ through conditional sampling: for any $t\in[T]$, let
@@
A larger value of $\ell_{\mathrm{PC}}(\bm{x})$ suggests that $\bm{x}$ is more likely to be closer to the model’s decision boundary. As we’ve explained in Section III-B, the prediction confidence of a neural network classifier can be an unreliable estimator. Therefore, we propose alternative guidance losses using more robust latent clustering-based approaches. Based on different clustering techniques, we again introduce two possible variants.
LCG-KM. For LCG-KM, we first obtain the latent embeddings of all the labeled data with respect to the intermediate model, namely $\{\bm{z}=h_{\hat{\theta}}(\bm{x}):\bm{x}\in\mathcal{S}_{l}\}$, and then apply k-means clustering algorithm to partition the embeddings into $k$ clusters. Note that this step is different from LCS-KM, because we now do not pre-generate the unlabeled dataset. Let $\{\bm{\mu}_{1},\bm{\mu}_{2},\ldots,\bm{\mu}_{k}\}$ denote the centroids corresponding to the $k$ returned clusters. We define the guidance loss as:
@@
where $\lambda\geq 0$ balances the trade-off between high-quality image generation captured by the original DDPM loss and boundary-seeking data generation driven by our guidance loss. By penalizing the model weights through the regularization term, our method gradually pushes $\theta$ towards higher probabilities of generating boundary-adjacent data using stochastic gradient descent (SGD). Our later experiments show that when starting from a pre-trained DDPM (i.e., on CIFAR-10), a few fine-tuning epochs are sufficient to bias the diffusion model to generate the desirable subset of synthetic images.
IV Experiments
In this section, we test the performance of our data reduction strategies proposed in Section III compared with baseline SSAT algorithms. In particular, we first conduct experiments to evaluate our selection-based methods in terms of robustness and efficiency advantages when integrated into the SSAT pipeline (Section IV-A). Then, we highlight the computational advantages of our approaches, particularly their capabilities of accelerating the convergence and reducing the total runtime of SSAT (Section IV-B). Finally, we study the effectiveness of our DDPM fine-tuning schemes in further improving the efficiency of SSAT algorithms under a generative setup (Section IV-C). To begin with, we lay out the main experimental setup. All our experiments are conducted using a single A100 GPU. All the remaining details, including all the hyperparameters, for reproducing our experiments are provided in Appendix B.
@@ IV Experiments
Dataset. We test two image benchmarks in our evaluation, SVHN [26] and CIFAR-10 [27], both widely used in prior adversarial ML literature. We consider two scenarios in terms of the source of the utilized unlabeled dataset $\mathcal{S}_{u}$ for testing our selection-based strategies: (i) external–using an external dataset, and (ii) generative–using a synthetic dataset generated by some pre-trained generative model. For external scenarios, we follow the implementation protocols outlined by Carmon et al. [11]. To be more specific, SVHN models are trained on 7373K labeled digit images from the original dataset, supplemented by 531531K extra unlabeled SVHN images. For CIFAR-10, the original dataset comprises 5050K labeled training images and 1010K labeled test images, where models trained using SSAT are augmented by an external unlabeled dataset $\mathcal{S}_{u}$ with 500500K images sampled from the 8080M Tiny Images (80M-TI) dataset. For generative setup, we adopt 11M synthetic images generated by DDPM for both datasets, following the protocols documented by Gowal et al. [13].
@@
Configuration. Following prior literature, we mainly consider $L_{\infty}$-norm bounded perturbations with $\epsilon=0.015$ on SVHN and $\epsilon=0.031$ on CIFAR-10. We adopt WideResNet (WRN) architectures and train models using TRADES [9] on both labeled and unlabeled data, coupled with pseudo labels returned by the intermediate model. In Appendix D, we evaluate the generalizability of our methods for varying $\ell_{\infty}$ perturbation size, vanilla PGD-based adversarial training with varying attack steps, and $\ell_{2}$ perturbations. To standardize training, an epoch is defined as processing 5050K data points, regardless of the total dataset size. Prior works also adopted this approach [11, 13], which ensures that training time per epoch remains consistent across datasets of varying sizes.
SSAT. For experiments with external unlabeled data, SSAT algorithms are trained for a total of $200$ epochs, whereas experiments with generated data extend to $400$ epochs. To compare the performance of our methods in a fair way, models are saved every $25$ epoch of SSAT, and the model achieving the highest robust accuracy is selected as the best model. This approach aligns with the early stopping practices commonly used in adversarial ML literature [9, 28]. In experiments with external data and ratio parameter $\alpha\in\{10\%,20\%\}$, we consistently observe peak robust accuracy around $100$ epochs for CIFAR-10 and $75$ epochs for SVHN. Conversely, when no data selection scheme is applied, peak performance is delayed to approximately $200$ epochs. Total training time is reported as the duration required to achieve the best model performance.
Intermediate Model. Our data selection and guided diffusion schemes require training an intermediate model $f_{\hat{\theta}}$, which we adopt the same architecture as the final model. This intermediate model is trained using standard supervised learning for $100$ epochs on a WRN-28-10 architecture, which takes around $55$ minutes. Notably, this time is excluded from the reported total runtime. The intermediate model fulfills two critical roles: it not only facilitates the pseudo-labeling of unlabeled data but also constitutes a necessary component for implementing any of our proposed strategies. As such, this training step is integral to the overall workflow, irrespective of whether a specific selection or fine-tuning method is employed.
Evaluation Metric. To test the robustness of the final model returned by SSAT, we mainly use multi-step PGD attacks (PGD) [7], due to their broad adoption in adversarial ML. We also report robust accuracies against AutoAttack (AA) [29] for more rigorous robustness evaluation and the model’s clean accuracy on normal inputs (Clean) as a reference.
@@
Dataset Ratio $\bm{\alpha}$ Method External Unlabeled DDPM-Generated
@@
IV-A Strategic Unlabeled Data Selection
SVHN. Our evaluation results on SVHN are summarized in Table I, which demonstrate the efficacy of the proposed data selection schemes when employed in SSAT algorithms trained using a WRN-16-8 architecture with varying data ratios $\alpha\in\{1\%,10\%,20\%\}$, with comparisons to baseline SSAT performance with $\alpha\in\{0\%,100\%\}$. Specifically, employing our LCS-KM strategy with $10\%$ or $20\%$ of extra unlabeled data achieves robust accuracies of $86.3\%$ and $86.6\%$, respectively, which matches and even slightly outperforms the $86.0\%$ robustness achieved using the entire unlabeled dataset. In contrast, random selection with $20\%$ of the data yields a robust accuracy of only $82.3\%$, underscoring the superior efficacy of informed selection strategies such as LCS-KM. The robust accuracies evaluated using AutoAttack (i.e., the AA column of Table I) demonstrate similar patterns compared to those using multi-step PGD Attacks (i.e., the PGD column of Table I). Additionally, alternative methods, including PCS and LCS-GMM, consistently outperform random selection, confirming the effectiveness of our insights. Notably, the robustness of models trained using SSAT peaks when the boundary ratio parameter $\beta$ is inversely proportional to the selection ratio $\alpha$, suggesting that smaller $\beta$ values are favorable as the data inclusion proportion increases. We provide detailed sensitivity analyses of these hyperparameters in Section V-C.
Our evaluations with 11M DDPM-generated data highlight that when using $10\%$ to $20\%$ of unlabeled data selected via LCS-KM, SSAT can achieve comparable robustness to the SSAT baseline. Since all SSAT algorithms listed under the column “DDPM-Generated” in Table I require the same amount of pre-generation time, we exclude it from the total runtime. In particular, training SSAT with reduced data requires only $75$ epochs to achieve the peak robust accuracy, while training with the full dataset requires up to $400$ epochs. Such a large reduction in computation time underscores the practical advantages of the proposed data selection strategies, which are further elaborated in Section IV-B. Besides, we observe in our experiments that using ground-truth labels reveals negligible differences compared to pseudo-labeled data, indicating that the primary performance gains stem from the selection strategy itself rather than label accuracy (see Table IV in Appendix C).
CIFAR-10. We adopt a WRN-28-10 architecture for CIFAR-10 experiments. To assess the generalizability of our findings, we also conduct ablations on ResNet-18, which are detailed in Table V in Appendix C. Across varying ratios of utilized unlabeled data $\alpha$ and different data selection schemes, Table I reveals that selecting $20\%$ of the external data using LCS-KM attains robust accuracy of $60.7\%$, comparable to the $62.5\%$ obtained with the entire dataset. In comparison, random selection with the same ratio parameter achieves only $57.8\%$, underscoring the advantages of our strategic LCS technique with k-means. Similar to SVHN experiments, we extend our analysis to CIFAR-10 with synthetically generated data. Using 11M DDPM-generated CIFAR-10-like images, we discover that using just $20\%$ of the generated data via LCS-KM results in performance nearly identical to that achieved with the full dataset. These trends are also consistent across experiments with $\ell_{2}$ perturbations (see Figure 10 for supporting evidence).
Our findings emphasize the robustness and adaptability of the proposed data selection schemes, especially for LCS-KM, even under varying experimental configurations. In terms of computational efficiency, CIFAR-10 experiments highlight the significant advantages of data selection. When leveraging selected external unlabeled data, optimal robust performance was achieved after $100$ epochs of training. By contrast, using the full unlabeled or generated datasets without selection required extended training durations of $200$ and $400$ epochs, respectively, to reach comparable performance. These results validate the efficacy of our latent clustering-based selection approach in enhancing semi-supervised adversarial training. The proposed method not only maintains the robust accuracy of the final model, $\theta_{\mathrm{final}}$, but also substantially reduces computational costs and accelerates the training process, underscoring its practical utility and scalability in SSAT frameworks.

### v3

@@ IV-A Prioritize Boundary-Adjacent Unlabeled Data
IV-B Strategic Selection via Latent Clustering
PCS. To identify a core subset of boundary-adjacent points while accounting for computational efficiency, we first propose a straightforward approach, coined prediction confidence-based selection (PCS), which makes use of the intermediate model $f_{\hat{\theta}}$ to compute a prediction confidence score for each unlabeled data point. Algorithm 1 in Appendix A presents the pseudocode of this selection scheme. Initially, all the unlabeled data points in $\mathcal{S}_{u}$ are sorted by their prediction confidence $\mathrm{Conf}(\cdot)$, with those exhibiting the lowest confidence scores being prioritized. The underlying assumption is that data points with low confidence scores are more likely to lie near the model’s decision boundary, making them ideal candidates for our selection. Note that the parameter $\beta\in[0,1]$ is introduced to balance the ratio between boundary and non-boundary points to avoid overfitting, which is used in all our proposed selection schemes. The reason for involving such a trade-off parameter will be further discussed in Section VI-C.
@@ IV-B Strategic Selection via Latent Clustering
In particular, we explore two classical clustering techniques for their unique benefits in identifying boundary points in the LCS framework. Algorithm 2 depicts the pseudocode of the methods. In particular, LCS with k-means (LCS-KM) generates latent representations of unlabeled data and clusters them based on Euclidean distances to the centroids, prioritizing data points that are equidistant from multiple centroids to effectively capture local geometric structures around decision boundaries. On the other hand, LCS with Gaussian mixture models (LCS-GMM) fits the latent representations to Gaussian mixture models, using posterior probabilities across multiple fitted Gaussians to identify points that are likely near decision boundaries. Compared to PCS, both LCS-KM and LCS-GMM leverage latent space clustering to provide a more accurate characterization of boundary vulnerabilities.
@@
where $\pi_{j}$ is the mixing coefficient for the $j$-th Gaussian component, and $\mathcal{N}(\bm{z}\mid\bm{\mu}_{j},\mathbf{\Sigma}_{j})$ denotes the Gaussian distribution with mean $\bm{\mu}_{j}$ and covariance $\Sigma_{j}$. We focus on data points with similar top posterior probabilities, suggesting that they are near the boundary between Gaussian distributions. Specifically, for each $\bm{z}$, we calculate the difference between the highest and the second-highest posterior probabilities: $\Delta p=|p_{1}(\bm{z})-p_{2}(\bm{z})|$, where $p_{1}(\bm{z})$ and $p_{2}(\bm{z})$ are the highest and second-highest posterior probabilities, respectively. Data points with the smallest $\Delta\gamma$ values are selected, as they are more likely to be near decision boundaries. Finally, the top $\alpha|\mathcal{S}_{u}|$ unlabeled data with the smallest $\Delta p$ values are used to form the selected subset. As will be illustrated in Section V, the robustness performance of SSAT can be largely maintained when using LCS-KM and LCS-GMM to select a small subset of unlabeled data (e.g., $\alpha=20\%$). More detailed discussions of the difference between the proposed data selection schemes are provided in Section VI-B, where we visualize the selected unlabeled data in a two-dimensional latent space.
IV-C Guided Diffusion via Latent Clustering
@@
While our selection strategy can effectively reduce the unlabeled data set size and lower the training time, it has a critical drawback: when SSAT employs a generative approach, all our previous methods require generating the full synthetic unlabeled dataset as a prerequisite before selection, which is inefficient in terms of both generation time and storage requirement. As shown in Table III, generating 1M CIFAR-10-like synthetic images via DDPM requires around $4$ hours on a single A100 GPU. Note that the best current SSAT method, as documented in the RobustBench leaderboard [8], utilizes a few hundred million synthetically generated images, suggesting that the total generation time can be a significant computational overhead. Therefore, a natural question arises: whether pre-generating synthetic data and then applying strategic selection schemes can be further improved.
In this section, we introduce an alternative data generative pipeline, detailed in Algorithm 3, which answers the above question affirmatively. In particular, we fine-tune a pre-trained DDPM using novel guidance losses to achieve more efficient unlabeled data generation, building on the similar insights presented in the previous sections. In this way, we directly generate a small, critical set of boundary-adjacent data points, thereby avoiding redundant data generation and further reducing the overall computation overhead.
@@
DDPM. Diffusion models represent a milestone in the history of generative modeling, achieving state-of-the-art performance for high-quality image generation and synthesis. A denoising diffusion probabilistic model (DDPM) is a deep neural network trained to learn how to step-by-step reverse the forward process of adding Gaussian noise to an image [21]. By gradually removing Gaussian noise through the reverse process, DDPM can thus generate real-looking images from pure noise. Let $q_{0}$ be a real image distribution over $\mathbb{R}^{d}$ and $\mathcal{N}(\bm{0},\mathbf{I})$ be the noise distribution. The forward process defines how to map images $\bm{x}_{0}\sim q_{0}$ to Gaussian noises $\bm{x}_{T}\sim\mathcal{N}(\bm{0},\mathbf{I})$ through conditional sampling: for any $t\in[T]$, let
@@
A larger value of $\ell_{\mathrm{PC}}(\bm{x})$ suggests that $\bm{x}$ is more likely to be closer to the model’s decision boundary. As we’ve explained in Section IV-B, the prediction confidence of a neural network classifier can be an unreliable estimator. Therefore, we propose alternative guidance losses using more robust latent clustering-based approaches. Based on different clustering techniques, we again introduce two possible variants.
LCG-KM. For LCG-KM, we first obtain the latent embeddings of all the labeled data with respect to the intermediate model, namely $\{\bm{z}=h_{\hat{\theta}}(\bm{x}):\bm{x}\in\mathcal{S}_{l}\}$, and then apply k-means clustering algorithm to partition the embeddings into $k$ clusters. Note that this step differs from LCS-KM, as we no longer pre-generate the unlabeled dataset. Formally, let $\{\bm{\mu}_{1},\bm{\mu}_{2},\ldots,\bm{\mu}_{k}\}$ denote the centroids corresponding to the $k$ returned clusters, then we define the guidance loss for LCG-KM as:
@@
where $\lambda\geq 0$ balances the trade-off between high-quality image generation captured by the original DDPM loss and boundary-seeking data generation driven by our guidance loss. By penalizing the model weights through the regularization term, our method gradually pushes $\theta$ towards higher probabilities of generating boundary-adjacent data using stochastic gradient descent (SGD). Our later experiments show that when starting from a pre-trained DDPM (i.e., on CIFAR-10), a few fine-tuning epochs are sufficient to guide the diffusion model to generate the desirable subset of synthetic images.
V Experiments
In this section, we evaluate the performance of our data selection methods (Section V-A), discuss their computational advantages compared to full SSAT (Section V-B), and study the effectiveness of our guided DDPM fine-tuning schemes (Section V-C). All our experiments are conducted on a single A100 GPU. Below, we briefly introduce the main experimental setup, where full details are deferred to Appendix B.
@@ V Experiments
Dataset. We consider two benchmark image datasets, SVHN [30] and CIFAR-10 [26], and two scenarios in terms of the source of the unlabeled dataset $\mathcal{S}_{u}$ for evaluating our methods: (i) external–using an external dataset, and (ii) generative–using a synthetic dataset generated by some pre-trained generative model. For external scenarios, we follow the implementation protocols outlined by Carmon et al. [5]. To be more specific, SVHN models are trained on 7373K labeled digit images from the original dataset, supplemented by 531531K extra unlabeled SVHN images. For CIFAR-10, the original dataset comprises 5050K labeled training images and 1010K labeled test images, where models trained using SSAT are augmented by an external unlabeled dataset $\mathcal{S}_{u}$ with 500500K images sampled from the 8080M Tiny Images (80M-TI) dataset. For the generative setup, we use 11M synthetic images generated by DDPM for both datasets, following the protocols in Gowal et al. [17].
@@
Configuration. We consider $\ell_{\infty}$ perturbations with $\epsilon=0.015$ on SVHN and $\epsilon=0.031$ on CIFAR-10. We adopt WideResNet (WRN) architectures and train models using TRADES [50] on both labeled and unlabeled data, incorporating pseudo labels generated by the intermediate model. In Appendix D, we evaluate the generalizability of our methods for varying $\ell_{\infty}$ perturbation size, vanilla PGD-based adversarial training with varying attack steps, and $\ell_{2}$ perturbations. To standardize training, an epoch is defined as processing 5050K data points, regardless of the total dataset size. Prior works also adopted this approach [5, 17], which ensures that training time per epoch remains consistent across datasets of varying sizes.
@@
Dataset $\bm{\alpha}$ Method External Unlabeled DDPM-Generated
@@
SSAT. With external data, we train SSAT algorithms for a total of $200$ epochs using stepwise learning rate scheduling, whereas training is extended to $400$ epochs under generative setups. To ensure fair comparisons, models are saved every $25$ epoch, and the model achieving the highest robust accuracy is selected as the best model. This approach aligns with the early stopping practices widely used in prior literature [50, 33]. When the ratio of selected unlabeled data $\alpha\in\{10\%,20\%\}$, we consistently observe peak robust accuracy around $100$ epochs for CIFAR-10 and $75$ epochs for SVHN. Conversely, when no data selection is applied, peak robustness is delayed to approximately $200$ epochs. The total training time is reported as the duration required to achieve optimal performance.
Intermediate Model. Our data selection and guided diffusion approaches require training an intermediate model $f_{\hat{\theta}}$, which adopts the same architecture as the final model. The intermediate model fulfills two critical roles: it not only facilitates the pseudo-labeling of unlabeled data but also constitutes a necessary component for implementing any of our data reduction strategies. To be more specific, we train the intermediate model $f_{\hat{\theta}}$ using standard supervised learning for $100$ epochs on WRN-28-10, which takes approximately $55$ minutes. We exclude this time from the reported total runtime, as this step is integral to the overall pipeline of any tested SSAT algorithm.
Evaluation Metric. To evaluate the robustness of the final model returned by SSAT, we mainly employ multi-step PGD attacks (PGD) [27], due to their broad adoption in adversarial ML. We also report robust accuracies against AutoAttack (AA) [9] to provide a more rigorous robustness evaluation, as well as the model’s clean accuracy on normal inputs (Clean).
V-A Strategic Unlabeled Data Selection
Table I documents our evaluation results on SVHN and CIFAR-10, showing the efficacy of our data reduction schemes when incorporated into semi-supervised adversarial training under varying data ratios $\alpha\in\{1\%,10\%,20\%\}$, compared with baseline TRADES ($\alpha=0\%$) and full SSAT ($\alpha=100\%$).